OpenStudy (anonymous):

Given the linear equation y = -2x – 4, find the y-coordinates of the points (0, ), (-1, ), and (2, ). Please show all of your work. Plot these points and graph the linear equation.

OpenStudy (amistre64):

y = -2(0)-4 ; (0,y) y = -2(-1)-4 ; (-1,y) y = -2(2)-4 ; (2,y)

OpenStudy (anonymous):

just insert the given x-coordinates as x in the equation: y1=-2*0 -4 =-4 =>p1(0,-4) y2=-2*-1-4 =-2 =>p2(-1,-2) y3=-2* 2-4 =-8 =>p3(2,-8)

OpenStudy (anonymous):

how do i graph this!

OpenStudy (anonymous):

as it is a line, just draw a 2d cartesian coordinate system. then draw the points at their coordinates and then draw a line going through all of them. thats the line with the equation y=-2x-4.